seanspence:

If arc CE = 40°, and m∠CDE = 30°, what is arc AB? 70° 100° 140° 20°

dude:

1 attachment
dude:

To solve this, use the same equation you would find to use to solve for the angle \(m\angle CDE=\dfrac{far~arc-near~arc}{2}\) or \(m\angle CDE=\dfrac{arc~AB-arc~CE}{2}\) We are solveing for AB (Let's substitute what we have) AB = x \(30=\dfrac{x-40}{2}\) I assume you can solve from here